An expensive watch is powered by a​ 3-volt lithium battery expected to last five years. Suppose the life of the battery has a standard deviation of 0.7 year and is normally distributed.

a. Determine the probability that the​ watch's battery will last longer than 5.2 years.

b. Calculate the probability that the​ watch's battery will last more than 4.45 years.

c. Compute the​ length-of-life value for which 5​% of the​ watch's batteries last longer.
Answer #1

5.2-5.0 a) P(x>5.2) =P* 0.7 =P(Z 0.29) =1-P(Z<0.29) =1-0.6141 =0.3859 (using z value table) 5.2-5.0 Using excel =1-NORMSDISTc) P(X <x)=0.05 P(< *539) = 0.05 Z(0.05)=-1.645 x-5.0 = -1.645 0.7 x = 5.0-1.645*0.7 = 3.8485 = 3.85 Length of life value for